### v1

@@ 4.2 Effects of irradiation on fracture energy
$w_{\rm p}\left(\mathrm{dpa},T\right)=G_{\rm c}\left(\mathrm{dpa},T\right)-2\gamma_{\rm s}\left(\mathrm{dpa},T\right),\quad 2\gamma_{\rm{sep}}\left(\mathrm{dpa},T\right)\leq 2\gamma_{\rm{s}}\left(\mathrm{dpa},T\right)\leq G_{c}\left(\mathrm{dpa},T\right).$ (6)
@@
The radiation-induced ductile-to-brittle transition can be characterized by a decrease in $w_{\rm p}(\mathrm{dpa},T)$ as the radiation dose (dpa) increases. Radiation-induced defects affect dislocation motion, altering the balance between plasticity and bond rupture during crack propagation (see Section 3.3). The following regimes can be identified based on the value of $w_{\rm p}(\mathrm{dpa},T)$:
@@ 4.3 Energy dissipation in irradiated materials
In order to evaluate the plastic work $w_{\rm p}(\mathrm{dpa},T)$ as defined in Eq. (6), the critical energy release rate $G_{\rm c}\!\left(\mathrm{dpa},0\rm K\right)$ should be determined at absolute zero temperature. To pragmatically approximate $G_{\rm c}\!\left(\mathrm{dpa},0\rm K\right)$, a finite simulation temperature of $T$=10K is employed.
@@
T=300K; sample irradiated to 0.07 dpa
T=10K; sample irradiated to 0.07 dpa
Figure 11: Equivalent shear strain fields calculated using the Green–Lagrangian strain tensor [37] in a selected cross-section of the samples irradiated to 0.07 dpa at 300K and 10K.
Fig. 11 compares the equivalent shear strain fields, calculated similarly to Fig. 6, in cross-sections of samples irradiated to 0.07 dpa at 300K and 10K. At 300K (top row), the strain field exhibits a more widespread distribution than 10K (bottom row), with deformation occurring throughout the sample and strain localizing particularly around the crack and radiation-induced defects. This reflects thermally activated dislocation motion, which enables plasticity and contributes to energy dissipation. In contrast, at 10K (bottom row), the strain is highly concentrated around the crack tip and radiation defects, with minimal deformation in the surrounding material. The suppression of strain outside these localized regions indicates restricted dislocation mobility and reduced plasticity, leading to more brittle fracture behavior. This comparison demonstrates that separation energy $\gamma_{\rm sep}$ and simulations at 10K can serve as a near-brittle reference state for approximating $G_{\rm c}$(dpa, 0K) in Eq. (6).
@@
300K   unirradiated sample                   dpa=0.008                   dpa=0.038                        dpa=0.152
Refer to caption Refer to caption Refer to caption Refer to caption
10K    unirradiated sample                   dpa=0.008                   dpa=0.038                        dpa=0.152
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 12: Stress–strain curves and energy evolution for different irradiation levels at 300 K and 10 K, showing internal energy $\Delta U$ (red), external work $W$ (blue), and stress–strain response (black).
At 300K, all samples exhibit discrete steps in their post-peak stress–strain curves, corresponding to crack arrest at the knots of entangled dislocations (compare Fig. 9 and Section 3). These steps are also evident in the internal energy curve, reflecting energy relocation (or redistribution) as it diverts (or shifts) from bond rupture to dislocation nucleation. The unirradiated sample displays a characteristic post-yield plateau (Fig. 12, top-left), indicative of homogeneous plastic zone development through unimpeded dislocation motion. Irradiation disrupts this plasticity: especially visible at 0.008 and 0.152 dpa, the stress drops abruptly after yielding, indicating defect-dominated fracture driven by crack-void interactions with limited plastic dissipation (compare Section 3).
@@
Figure 13: Heat taken from the system $(-Q=W-\Delta U)$ as a function of projected crack surface area $A$ for a wide dpa range at (a) 300K and (b) 10K.
Fig. 13 shows the heat taken from the system ($-Q=W-\Delta U$) as a function of the projected crack surface area $A$ for all samples considered. Here, $W$ and $\Delta U$ correspond to those shown in Fig. 12, while the surface area $A$ represents the two-dimensional geometric projection of the crack’s surface onto the x–z plane. Note that the initial drop of $-Q$ in Fig. 13(a) corresponds to the elastic deformation stage at which the crack surface area remains unchanged at $A=40$ nm2, while the heat is supplied to the system ($Q>0$), as discussed above with reference to Fig. 12.
By following the approach of Xu and Demkowicz [20], the dependence of $-Q$ on $A$ can now be used to determine the critical energy release rate $G_{\rm c}$. Indeed, in the quasi-steady state crack propagation regime, the heat taken from the system by the thermostat corresponds to the energy dissipated as a result of crack propagation and dislocation activity, thus $\mathrm{d}E_{\rm dis}=-\mathrm{d}Q$. For each sample, $G_{\rm c}$ can thus be determined as the slope of the corresponding curve in Fig. 13 according to Eq. (2). Note that the initial stage (marked in blue) and the final stage (marked in purple) are excluded from calculations. The transient initial stage corresponds to the build-up of the plastic zone ahead of the crack tip. The final stage is omitted because of the direct interaction of the crack front with the boundary of the simulation box.
@@
The results show a trend consistent with an irradiation-induced ductile-to-brittle transition. At T=300K, $G_{\rm c}$ decreases over the examined dpa range, indicating suppressed energy dissipation mechanisms as defect density increases. The higher $G_{\rm c}$ at 300K is attributed to thermally activated processes that allow some dislocation-mediated energy absorption. At T=10K, $G_{\rm c}$ moves closer to the theoretical limit governed by intrinsic surface energy $\gamma_{\rm{sep}}$ (Fig. 14(b)), as the lack of thermal energy prevents significant dislocation motion, making bond rupture the predominant fracture mechanism.
@@
5 Summary and discussion
This study provides a comprehensive investigation into the fracture behavior of irradiated Fe–Ni–Cr alloys, revealing how radiation-induced defects govern the transition from ductile to brittle failure. Although the system investigated in this study does not exhibit fully brittle failure due to the limited defect density inherent to the small MD sample size, the observed embrittlement trends offer valuable insights into the underlying mechanisms governing this transition.
The crack propagation behavior in irradiated materials is governed by a complex interplay of mechanisms, as detailed in Section 3.3. Voids act as weak zones that accelerate fracture through coalescence, reducing the cohesive energy and providing pathways for crack growth. Dislocation loops obstruct dislocation motion, redirecting cracks and promoting branching, which alters the crack path and increases fracture surface area. At higher irradiation levels, amorphous zones form at crack tips, suppressing plasticity and promoting a more brittle failure mode. High defect densities also lead to crack branching and the initiation of secondary cracks, which coalesce with the main crack, resulting in irregular fracture patterns. Additionally, radiation-induced defects generate localized stress concentrations, further influencing crack propagation and fracture morphology. These mechanisms collectively contribute to embrittlement, with the balance between plasticity and bond rupture shifting as irradiation damage accumulates.
@@ 5 Summary and discussion
The study advances the field through three key contributions:
  • Defect-driven fracture pathways: By categorizing crack-void and crack-loop interactions, we establish that voids accelerate fracture by reducing cohesive strength, while loops impede propagation by pinning dislocations. This duality explains the transition from ductile tearing to brittle cleavage as dpa increases.
  • Microstructural-mechanical linkage: The correlation between defect density, crack morphology, and energy dissipation bridges atomic-scale simulations to macroscopic fracture resistance, offering a unified model for irradiation embrittlement.
  • Energy-based embrittlement criterion: Building on Xu and Demkowicz [20], we adapt the decomposition of $G_{\rm c}$ into brittle ($\gamma_{\rm s}$) and plastic ($w_{\rm p}$) contributions for irradiated materials. This framework provides a step toward quantifying how irradiation affects the fracture energy balance.
  • Quantitative description of irradiation-induced DBT: The study demonstrates a decrease in both $G_{\rm c}$ and $w_{\rm p}$ with increasing dpa, providing a quantitative manifestation of irradiation-induced ductile-to-brittle transition. While the observed embrittlement is relatively small due to the limited dpa range in our MD simulations, real-world scenarios with higher dpa values and greater defect densities are expected to exhibit more pronounced DBT effects.

### v2

@@ 4.2 Effects of radiation on fracture energy
$w_{\rm p}\left(\mathrm{dpa},T\right)=G_{\rm c}\left(\mathrm{dpa},T\right)-2\gamma_{\rm s}\left(\mathrm{dpa},T\right),\quad 2\gamma_{\rm{sep}}\left(\mathrm{dpa},T\right)\leq 2\gamma_{\rm{s}}\left(\mathrm{dpa},T\right)\leq G_{\rm c}\left(\mathrm{dpa},0{\rm K}\right).$ (6)
@@
The radiation-induced DBT can be characterized by a decrease in $w_{\rm p}(\mathrm{dpa},T)$ as the radiation dose (dpa) increases. Radiation-induced defects affect dislocation motion, altering the balance between plasticity and bond rupture during crack propagation (see Section 3.3). The following regimes can be identified based on the value of $w_{\rm p}(\mathrm{dpa},T)$:
@@ 4.3 Energy dissipation in irradiated materials
In order to evaluate the plastic work $w_{\rm p}(\mathrm{dpa},T)$ as defined in Eq. (6), the critical energy release rate $G_{\rm c}\!\left(\mathrm{dpa},0\rm K\right)$ should be determined at absolute zero temperature. To pragmatically approximate $G_{\rm c}\!\left(\mathrm{dpa},0\rm K\right)$, a finite simulation temperature of $T$=10K is employed,
@@
Fig. 8 compares the equivalent shear strain fields in cross-sections of samples irradiated to 0.07 dpa at 300K and 10K. At 300K (top row), the strain field exhibits a more widespread distribution than at 10K (bottom row), with deformation occurring throughout the sample and strain localizing particularly around the crack and radiation-induced defects. This reflects thermally activated dislocation motion, which enables plasticity and contributes to energy dissipation. In contrast, at 10K (bottom row), the strain is highly concentrated around the crack tip and radiation defects, with minimal deformation in the surrounding material. The suppression of strain outside these localized regions indicates restricted dislocation mobility and reduced plasticity, leading to more brittle fracture behavior. This comparison demonstrates that separation energy $\gamma_{\rm sep}$ and simulations at 10K can serve as a near-brittle reference state for approximating $G_{\rm c}$(dpa, 0K) in Eq. (6).
@@
Figure 8: Evolution of the equivalent shear strain (von Mises invariant of the Green–Lagrangian strain tensor [44]) in a selected cross-section of the samples irradiated to 0.07 dpa at 300K (top) and 10K (bottom).
Figure 9: Stress–strain curves and energy evolution for different irradiation levels at 300 K (top) and 10 K (bottom), showing internal energy $\Delta U$ (red), external work $W$ (blue), and stress–strain response (black).
Figure 10: Heat taken from the system $(-Q=W-\Delta U)$ versus projected crack surface area $A$ for various dpa levels at (a) 300 K and (b) 10 K. Blue and pink shaded regions denote excluded intervals corresponding to initial plastic zone development and final boundary effects, respectively. The initial drop in $-Q$ observed at 300 K (but absent at 10 K) corresponds to the elastic deformation stage.
At 300K, all samples exhibit discrete steps in their post-peak stress–strain curves, corresponding to crack arrest at the knots of entangled dislocations (compare Fig. 6 and Section 3). These steps are also evident in the internal energy curve, reflecting energy relocation (or redistribution) as it diverts (or shifts) from bond rupture to dislocation nucleation. The unirradiated sample displays a characteristic post-yield plateau (Fig. 9, top-left), indicative of homogeneous plastic zone development through unimpeded dislocation motion. Irradiation disrupts this plasticity: especially visible at 0.008 dpa and 0.152 dpa, the stress drops abruptly after yielding, indicating defect-dominated fracture driven by crack-void interactions with limited plastic dissipation (compare Section 3).
@@
Fig. 10 shows the heat taken from the system ($-Q=W-\Delta U$) as a function of the projected crack surface area $A$ for all samples considered. Here, $W$ and $\Delta U$ correspond to those shown in Fig. 9, while the surface area $A$ represents the two-dimensional geometric projection of the crack’s surface onto the $x$$z$ plane. Note that the initial drop of $-Q$ in Fig. 10(a) (observed at 300 K, but not at 10 K) corresponds to the elastic deformation stage at which the crack surface area remains unchanged at $A=40$ nm2, while the heat is supplied to the system ($Q>0$), as discussed above with reference to Fig. 9.
By following the approach of Xu and Demkowicz [21], the dependence of $-Q$ on $A$ can now be used to determine the critical energy release rate $G_{\rm c}$. Indeed, in the quasi-steady state crack propagation regime, the heat taken from the system by the thermostat corresponds to the energy dissipated as a result of crack propagation and dislocation activity, thus $\mathrm{d}E_{\rm dis}=-\mathrm{d}Q$. For each sample, $G_{\rm c}$ can thus be determined as the slope of the corresponding curve in Fig. 10 according to Eq. (2). Note that the initial stage (marked in blue) and the final stage (marked in pink) are excluded from calculations. The transient initial stage corresponds to the build-up of the plastic zone ahead of the crack tip. The final stage is omitted because of the direct interaction of the crack front with the boundary of the simulation box.
@@
The results show a trend consistent with the irradiation-induced DBT. At T=300K, $G_{\rm c}$ decreases over the examined dpa range, indicating suppressed energy dissipation mechanisms as defect density increases. The higher $G_{\rm c}$ at 300K is attributed to thermally activated processes that allow some dislocation-mediated energy absorption. At T=10K, $G_{\rm c}$ moves closer to the theoretical limit governed by intrinsic surface energy $\gamma_{\rm{sep}}$ (Fig. 11(b)), as the lack of thermal energy prevents significant dislocation motion, making bond rupture the predominant fracture mechanism.
@@
5 Discussion
This study provides a comprehensive investigation into the fracture behavior of irradiated Fe–Ni–Cr alloys, revealing how radiation-induced defects govern the transition from ductile to brittle failure. The simulated defect configuration reflects the damage accumulation characteristic of neutron irradiation, allowing for a direct examination of how radiation-induced defects distributed throughout the material influence crack propagation and reduce the capacity for plastic deformation [62, 63]. Although the system investigated in this study does not exhibit fully brittle failure due to the limited defect density inherent to the small MD sample size, the observed embrittlement trends offer valuable insights into the underlying mechanisms governing this transition.
The crack propagation behavior in irradiated materials is governed by a complex interplay of several mechanisms, as detailed in Section 3.3. Voids act as weak zones that accelerate fracture through coalescence, reducing the cohesive energy and providing pathways for crack growth. Dislocation loops obstruct dislocation motion, redirecting cracks and promoting branching, which alters the crack path and increases fracture surface area. At higher irradiation levels, amorphous zones form at crack tips, suppressing plasticity and promoting a more brittle failure mode. High defect densities also lead to crack branching and the initiation of secondary cracks, which coalesce with the main crack, resulting in irregular fracture patterns. Additionally, radiation-induced defects generate localized stress concentrations, further influencing crack propagation and fracture morphology. These mechanisms collectively contribute to embrittlement, with the balance between plasticity and bond rupture shifting as irradiation damage accumulates.
@@ 5 Discussion
The study advances the field through four key contributions:
  • Defect-driven fracture pathways:  By categorizing crack–void and crack–loop interactions, we establish that voids accelerate fracture by reducing cohesive strength, while loops impede propagation by pinning dislocations. This duality explains the transition from ductile tearing to brittle cleavage as dpa increases.
  • Microstructural–mechanical linkage:  The correlation between defect density, crack morphology, and energy dissipation bridges atomic-scale simulations to macroscopic fracture resistance, offering a unified model for irradiation embrittlement.
  • Energy-based embrittlement criterion:  Building on Xu and Demkowicz [21], we adapt the decomposition of $G_{\rm c}$ into brittle ($\gamma_{\rm s}$) and plastic ($w_{\rm p}$) contributions for irradiated materials. This framework provides a step towards quantifying how irradiation affects the fracture energy balance.
  • Quantitative description of irradiation-induced DBT:  The study demonstrates a decrease in both $G_{\rm c}$ and $w_{\rm p}$ with increasing dpa, providing a quantitative manifestation of the irradiation-induced DBT. While the observed embrittlement is relatively small due to the limited dpa range in our MD simulations, real-world scenarios with higher dpa values and greater defect densities are expected to exhibit more pronounced DBT effects.
These findings pertain specifically to austenitic Fe–Ni–Cr alloys with an fcc lattice, where defect distributions are largely isotropic due to uniform threshold displacement energies and minimal channeling effects. Ferritic-martensitic steels, also used in nuclear fusion applications, are characterized by a bcc structure and low nickel content to mitigate transmutation reactions and reduce helium accumulation. These differences may influence the overall crack behavior because the variations in microstructure and defect evolution may modulate the fracture resistance. Nonetheless, the core mechanisms driving irradiation-induced embrittlement, dislocation pinning, void coalescence, and crack branching are expected to be fundamentally similar in both fcc and bcc steels. Thus, while this work directly addresses fcc austenitic steels, its insights contribute to a broader understanding of irradiation embrittlement.
6 Conclusions
This work has systematically investigated the fracture mechanisms in irradiated alloys, emphasizing the crucial role of radiation-induced defects in embrittlement. Based on MD simulations, the analysis shows that irradiation decreases the material’s ability to accommodate plastic deformation by impeding dislocation motion and promoting crack growth through the absorption of radiation-induced voids. These voids act as weak zones that facilitate crack propagation, collectively causing a transition from ductile to more brittle fracture behavior.
Our energy decomposition framework determines embrittlement by correlating atomistic defect-induced changes in fracture energy components with the macroscopic loss of plastic dissipation capacity. By separating total fracture energy into surface energy (crack surface creation) and plastic work (dislocation-mediated inelastic deformation), it provides a robust parameter that translates atomic-scale damage effects into continuum-scale fracture behavior.
@@ Acknowledgments
Appendix A Evolution of defect densities in irradiated samples
Supplementary data provided in this appendix illustrate the evolution of radiation-induced defect densities in the irradiated samples discussed in Section 2.1.
Figure A.1 shows how the densities of radiation-induced dislocation loops and voids evolve with increasing irradiation dose. The density of radiation defects increases gradually at lower dpa, reflecting the early stages of dislocation loop nucleation and the initial aggregation of vacancies into stable void structures. The absence of perfect dislocations at dpa = 0.008 reflects the early stage of defect accumulation, where primarily Shockley partials and small defect clusters are present. Perfect dislocations tend to form at higher dpa levels as overlapping cascades drive defect evolution. Void density includes all voids regardless of size. At higher dpa levels, the accumulation and interaction of dislocations lead to the formation of various dislocation types and a more complex defect network.
Figure A.1: Density of dislocation loops (a) and voids (b) as a function of the irradiation dose in dpa obtained from MD simulations of collision cascades.
Appendix B Effect of boundary conditions on fracture behavior
Figure B.1: Effect of boundary conditions on crack propagation. Top-left panel A show two variations of the ssp configuration: (i) with fixed lateral boundaries (left), corresponding to the setup used in the main study, and (ii) with free lateral boundaries (right). Bottom-left panel B shows the ffp configuration: (i) with constrained lateral edges and a vacuum region (light blue), and (ii) with unconstrained sides leading to pronounced necking. Panel C on the right shows the ppp setup with periodic boundary conditions for two crack locations: at the sample edge (left) and at the center (right). These conditions lead to periodic effects such as mirrored crack propagation (purple) and structural duplication (blue).
To assess how boundary constraints affect the physical accuracy of crack propagation in atomistic simulations, we performed a comparative analysis of three representative boundary setups commonly used in MD studies. Specifically, we examined: (i) shrink-wrapped lateral boundaries with fixed displacements applied at the top and bottom (ssp), (ii) fixed boundaries in all directions with vacuum padding (ffp), and (iii) fully periodic boundary conditions (ppp). All simulations were conducted under uniaxial tensile loading. Representative snapshots of the resulting deformation patterns and fracture evolution are provided in Fig. B.1. Such boundary configurations have also been investigated in previous studies focusing on fracture behavior, including fully periodic (ppp) [64], fully free (ffp), and non-periodic shrink-wrapped (ssp) systems [23, 21, 65, 8].
The upper-left panel A of Fig. B.1 illustrates two variations of the ssp configuration. The left snapshot in panel A corresponds to the ssp setup employed throughout our study, where atoms at the lateral edges (in the $x$-direction) are constrained to prevent transverse motion. In contrast, the right snapshot in panel A depicts a similar ssp configuration without lateral constraints, resulting in pronounced necking near the edges. The released lateral strain causes localized narrowing, which impedes the crack front and obstructs its advancement.
The ssp setup imitates a tensile test with clamped top and bottom edges, closely resembling experimental conditions where deformation is restricted by grips. The applied displacement is transmitted through fixed atomic layers at the top and bottom of the sample, inducing a deformation that gradually propagates through the lattice.
In contrast, the ffp setup (lower-left panel B) enforces affine deformation by globally modifying the simulation box dimensions. While this method is efficient and stable, it imposes a homogeneous strain field that accelerates local crack opening. The applied global strain leads to earlier failure initiation and can obscure important mechanisms such as strain localization or dislocation emission. This approach can approximate bulk material behavior, and the imposed uniform deformation limits local surface-related effects.
The lower-left panel B of Fig. B.1 illustrates two variants of the ffp boundary condition. The left snapshot in panel B corresponds to the ffp setup where atoms at the lateral edges (in the $x$-direction) are constrained to prevent transverse motion. To prevent atom deletion at the sample edges, a vacuum region (light blue) is introduced. This setup enforces affine deformation by globally modifying the simulation box dimensions. While efficient and stable, it imposes a homogeneous strain field that accelerates local crack opening and leads to necking near the constrained edges (compare snapshots for the same time steps in panels A and B). The applied global strain causes earlier failure initiation and can obscure important mechanisms such as strain localization or dislocation emission. In contrast, the right snapshot in panel B shows a similar ffp configuration without lateral constraints, leading to more freedom at the edges and resulting in more pronounced necking and different deformation behavior. This approach can approximate bulk material behavior. The imposed uniform deformation limits local surface-related effects and does not fully replicate free surface conditions essential for fracture studies.
The large panel on the right side of Fig. B.1 corresponds to the ppp configuration. Two cases are shown: with the crack initially placed at the left edge (top left) and at the center of the sample (top right). Due to periodic boundary conditions, mirrored cracks appear on the opposite side (highlighted in purple), and structural features are periodically duplicated (marked in blue). These effects modify the deformation characteristics and should be considered in the overall assessment of the system’s behavior.
The constrained boundary ssp setup used in this work has been selected as a compromise that preserves realistic plasticity and dislocation behavior while preventing unphysical deformation modes.
Appendix C Strain field maps of irradiated and unirradiated samples
Figure C.1: Evolution of the equivalent shear strain (von Mises invariant of the Green–Lagrangian strain tensor [44]) in a selected cross-section of unirradiated (top) and irradiated (bottom) samples.
This appendix presents supplementary strain field maps used to visualize deformation behavior in the irradiated and unirradiated samples discussed in Section 2.2. Fig. C.1 presents equivalent shear strain fields in both unirradiated and irradiated samples across various time steps. The equivalent shear strain is calculated by OVITO as the von Mises invariant of the Green–Lagrangian strain tensor [44]. The images show clear differences in strain localization patterns, where the irradiated samples exhibit concentrated strain fields at radiation-induced defects.